Activity-Based Product Costing Roberts Company produces two weed eaters: basic and advanced. The company has four activities: machining, engineering, receiving, and inspection. Information on these activities and their drivers is given below. Basic Advanced Total Units produced 130,000 390,000 — Prime costs $9,880,000 $42,120,000 $52,000,000 Machine hours 130,000 650,000 780,000 Engineering hours 500 4,500 5,000 Receiving orders 500 1,500 2,000 Inspection hours 1,000 2,000 3,000 Overhead costs: Machining $8,580,000 Engineering 3,000,000 Receiving 440,000 Inspecting products 360,000 Required: 1. Calculate the four activity rates. Machining rate $ per machine hour Engineering rate $ per hour Receiving rate $ per order Inspecting rate $ per hour 2. Calculate the unit costs using activity rates. Round your answers to the nearest cent. Unit cost Basic $ per unit Advanced $ per unit Calculate the overhead cost per unit. Round your answers to the nearest cent. Overhead Cost Basic $ per unit Advanced $ per unit 3. If consumption ratios instead of activity rates were used to assign costs instead of activity rates, show the cost assignment for the inspection activity. Cost assignment Basic $ Advanced $
Answer #1

1.

Machining rate = $8,580,000/780,000 = $11 per machine hour

Engineering rate = $3,000,000/5,000 = $600 per engineering hour

Receiving rate = $440,000/2,000 = $220 per inspecting hour

Inspecting rate = $360,000/3,000 = $120 per inspection hour

2.

Basic Advanced
Prime cost $9,880,000 $42,120,000
Overhead costs:
Machining 1,430,000 (130,000*$11) 7,150,000 (650,000*$11)
Engineering 300,000 (500*$600) 2,700,000 (4,500*$600)
Receiving 110,000 (500*$220) 330,000 (1,500*$220)
Inspecting 120,000 (1,000*$120) 240,000 (2,000*$120)
Total overhead cost 1,960,000 10,420,000
Total cost $11,840,000 $52,540,000
Units produced 130,000 390,000
Unit cost $91.08 $134.72
Overhead cost per unit $15.08 $26.72

3.

Cost assignment:

Basic = $360,000*1,000/3,000 = $120,000

Advanced = $360,000*2,000/3,000 = $240,000